OpenStudy (anonymous):

A rock is launched from a cannon. Its height, h(x), can be represented by a quadratic function in terms of time, x, in seconds. After 1 second, the rock is 58 feet in the air; after 2 seconds the rock is 112 feet in the air. Find the height, in feet, of the rock after 10 seconds in the air.

OpenStudy (anonymous):

multiple choice?

OpenStudy (anonymous):

... h(1) = a - b = 58 => a = b + 58 h(2) = 2a - 4b = 112 2(b + 58) - 4b = 112 - 2b + 116 = 112 2b = 4 => b = 2 so a = 2 + 58 = 60 so h(x) = 60x - 2x^2 so h(10) = 600 - 200 = ?

OpenStudy (anonymous):

no its not multiple choice

OpenStudy (anonymous):

From what i just put can you find the answer

OpenStudy (anonymous):

so h(10)= 600- 200 = 400 FT